OpenStudy (anonymous):

making waffles for 3 people. 2/3 cup of bleach flour and 1/4 cup of wheat flour how many cups of wheat flour does been need ? means to multiply add or subtract ?

OpenStudy (dumbcow):

Im guessing thats a recipe for 1 waffle or 1 serving so multiply by 3 to serve people

OpenStudy (anonymous):

1/4,2/3=4/12

OpenStudy (dumbcow):

?? 3*1/4 = 3/4

OpenStudy (dumbcow):

maybe i don't understand the problem what is it they are asking?

OpenStudy (dumbcow):

if you're adding them, 1/4 + 2/3 = 3/12 + 8/12 = 11/12